In triangle ABC, median AD is perpendicular to median BE. Find AB if BC=6 and AC=8
1
Expert's answer
2019-12-11T10:43:31-0500

Let O be the intersection point of the medians. Medians are divided in the ratio 2: 1 at the intersection, according to the property of medians.


Let Y = OD, X = OE.


Then OB = 2X, OA = 2Y.


Compose the system:


Y2 + 4X2 = 9

X2 + 4Y2 = 16


Express X = "\sqrt{16-4Y^2}" ;


Then

Y2 = 9-4 * (16-4Y2) = 9-64 + 16Y2

15Y2 = 55


"Y=\sqrt{\frac{11}{3}}, \," "X=\sqrt{16-4 \cdot \frac{11}{3}}=\sqrt{\frac{4}{3}}"


For side AB we get the equation:


AB2 = 4 (X2 + Y2) = 4 * (4/3 + 11/3) = 4 * 15/3 = 20


"AB=\sqrt{20}=2\sqrt{5}" . Answer: "2\sqrt{5}."